Find the Reynolds number if a fluid of viscosity 0.4 Ns/m2 and relative density of 900 kg/m3
through a 20 mm pipe with a velocity of 2.5 m/s. Is the flow Laminar or Turbulent?
Note: If Reynolds < 2000 the flow is laminar
If Reynolds > 4000 the flow is turbulent
If 2000 < Re < 4000 it is transitional flow
Re =
1
Expert's answer
2020-06-04T09:58:08-0400

The Reynolds number is given by the following expression:


"\mathrm{Re} = \frac{\rho u D}{\mu}"

where "\rho = 900 kg/m^3" is the relative density of the fluid, "u = 2.5 m\/s" is the velocity of the fluid, "D = 0.02 m" is the diameter of the pipe and "\mu = 0.4 N\cdot s/m^2".

Substituting values, obtain:


"\mathrm{Re} = \frac{900\cdot 2.5 \cdot 0.
02}{0.4} = 112.5"

As far as 112.5<2000, the flow is laminar.


Answer. Re = 112.5, laminar flow.